OpenStudy (anonymous):

Medal and fan reward I KNOW HOW TO WORK THIS OUT, I JUST WANT THE ANSWER (i'm just double checking to see if I'm right or not) a) If a sample containing 2.00 mL of nitroglycerin (density= 1.592 g/ml) is detonated, how mahy total moles of gas are produced? b) If each mole of gas occupies 55 L under the conditions of the explosion, how many liters of gas are produced? c) How many grams of N2 are produced in the detonation?

OpenStudy (anonymous):

What are you answers?

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

a) .102 b) 5.59 c) .437 (a friend of sort of helped me with it)

OpenStudy (anonymous):

@Shalante what u get?

OpenStudy (anonymous):

Yes that is correct!

OpenStudy (anonymous):

wait PLEASE @Shalante can you PLEASE explain to me how you got letter c

OpenStudy (anonymous):

Yes Grams of nitroglycerin is 3.184grams Molar mass of nitroglycerin is 227g/mol There are 6 moles of N2 for every 4 mols of nitroglycerin (C3H5N3O9) N2 molar mass is 28g/mol

OpenStudy (anonymous):

\[\frac{ 3.184g C _{3}H _{5}N _{3}O _{9} }{ 1} \times \frac{ 1mol C3H5N3O9 }{ 227g C3H5N3O9 } \times \frac{ 6mol N2 }{ 4mol C3H5N3O9 } \times \frac{ 28.00g N2 }{ 1mol N2 }\]

OpenStudy (anonymous):

when I did that I got .58

OpenStudy (anonymous):

Detonation of nitroglycerin reaction \[4C _{3}H _{5}N _{3}O _{9}\rightarrow 12CO _{2}+ 6N _{2}+O _{2}+10H _{2}O\]

OpenStudy (anonymous):

Lemme see.

OpenStudy (anonymous):

but when I multiplied I got .589 @ Shalante what am I doing wrong

OpenStudy (anonymous):

@Shalante

OpenStudy (anonymous):

Gimme a second.

OpenStudy (anonymous):

K THANKS! YOU'RE THE BEST

OpenStudy (anonymous):

Actually it is 0.58. I am right.

OpenStudy (anonymous):

Your friend was wrong.

OpenStudy (anonymous):

oh ok then

OpenStudy (anonymous):

I trust you

OpenStudy (anonymous):

Haha, I am glad you told me to show you the steps.

OpenStudy (anonymous):

The dimensional analysis looked close between 0.30-0.7 so I assumed you got it right. Thats is actually better for me to do it than check it lol.

OpenStudy (anonymous):

lol and for b it's just 55L times the answer I got for a right?

OpenStudy (anonymous):

Yeah because they to want to know the number of moles with a different amount of liter.

OpenStudy (anonymous):

all right thx

OpenStudy (anonymous):

No problem! I need to sleep now. Good luck with the rest